OpenStudy (anonymous):

five times a number k inus 9 is two thirds of a number j

OpenStudy (gorv):

well let number =x

OpenStudy (gorv):

5*k-9=j*2/3

OpenStudy (anonymous):

ok thnx

OpenStudy (gorv):

wlcm

OpenStudy (anonymous):

how would I graph [y+3]=-2 as an inequality\[\left| ? \right|\]

OpenStudy (gorv):

[y+3] is it like this or |y+3| ??